Circular No. 5600

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET HELIN-LAWRENCE (1992q)

Further precise positions have been reported as follows:

1992 UT α_{2000} δ_{2000} m_{1} Observer
Aug. 31.34948 2^{h}02^{m}49.^{s}74 −13^{o}22′08′′.7 15 Helin
31.44045 2 02 46.22 −13 24 54.9  _{”}
31.69462 2 02 37.19 −13 32 49.2 15 Kojima
31.70978 2 02 36.60 −13 33 17.3  _{”}
31.72118 2 02 36.05 −13 33 37.9 15 Seki
31.73652 2 02 35.55 −13 34 07.3  Kojima

E. F. Helin, K. J. Lawrence, J. Alu and P. Rose (Palomar). 0.46-m Schmidt.

Measurer Rose. T. Kojima (YGCO Chiyoda Station). Communicated by S. Nakano. T. Seki (Geisei). Diffuse, strong condensation. Communicated by Nakano.

The following parabolic orbital elements, from nine observations Aug. 29-31, are better determined than those of comet 1992p:

T = 1993 Mar. 13.416 TT  ω = 270.^{o}795
Ω = 194.829 }2000.0
q = 1.96951 AU  i = 107.405
1992 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 26 2^{h}05.^{m}49 −10^{o}43.^{′}9 2.362 3.053 124.^{o}7 15.^{o}8 15.2
31 2 03.03 −13 11.3 2.265 3.012 129.6 15.0 15.1
Sept. 5 1 59.71 −15 51.1 2.177 2.970 134.3 14.1 14.9
10 1 55.46 −18 42.4 2.099 2.930 138.4 13.2 14.8
15 1 50.19 −21 43.5 2.032 2.889 141.7 12.5 14.6
20 1 43.86 −24 52.0 1.977 2.849 143.7 12.0 14.5
25 1 36.39 −28 04.5 1.934 2.809 144.1 12.1 14.4
30 1 27.79 −31 16.8 1.904 2.769 142.9 12.6 14.3
Oct. 5 1 18.10 −34 24.4 1.888 2.730 140.0 13.6 14.2

CATALOGUE OF COMETARY ORBITS 1992 Further to IAUC 5598, checks for the MS-DOS diskette should be made

out to “Central Bureau for Astronomical Telegrams”. Either 5.25-inch and 3.5-inch diskettes are available for the same amount of $100.00.

1992 September 1 Brian G. Marsden